# Structured Packing in LLM Training Improves Long Context Utilization

### v8

@@ Abstract
Recent advancements in long-context large language models have attracted significant attention, yet their practical applications often suffer from suboptimal context utilization. This study investigates structuring training data to enhance semantic interdependence, demonstrating that this approach effectively improves context utilization. To this end, we introduce the Structured Packing for Long Context (SPLiCe) method, which utilizes retrieval to collate mutually relevant documents into long and coherent training examples. We validate SPLiCe empirically across models of varying sizes — $3$B, $7$B, and $13$B — achieving improved performance in long-context tasks, such as Qasper and HotpotQA. Remarkably, even brief fine-tuning with SPLiCe is sufficient to realize these benefits. Additionally, SPLiCe effectively mitigates the ‘lost-in-middle’ phenomenon often observed in large models. Our comprehensive analysis of SPLiCe explores its design choices and reveals intriguing transfer effects; for instance, training on programming code enhances performance on natural language tasks. 11footnotetext: University of Warsaw, 2Ideas NCBR, 3University of Edinburgh, 4Google DeepMind, 5Institute of Mathematics, Polish Academy of Sciences, 6deepsense.ai
Accepted to AAAI’25
@@ 1 Introduction
Large language models (LLMs) (Brown et al., 2020; Chowdhery et al., 2022; Lewkowycz et al., 2022; OpenAI, 2023b; Bai et al., 2023) have transformed the field of AI. Recently, the field has observed the rise of long context Language Models (LCLMs) that promise to unveil novel and powerful capabilities (Anthropic, 2023; OpenAI, 2023a; Gemini Team, 2024). However, their ability to process long context is not always as effective as one hopes for. Indeed, several studies have highlighted an important limitation: when processing prompts composed of multiple documents, LCLMs frequently encounter difficulties in accurately extracting relevant information (Tworkowski et al., 2023; Liu et al., 2023; Shi et al., 2023a). Additionally, they typically find it challenging to utilize information from the middle of their inputs (Liu et al., 2023), even on simple synthetic retrieval tasks (Li et al., 2023a). Understanding these issues is vital for advancements in LCLM technologies and calls for systematic research.
@@
In this work, we take a step towards better context utilization in LCLMs. We focus on training data, keeping other components, such as the architecture and training objectives, unchanged. The broad question is how to organize training data to enhance long context capabilities? Such perspective has received some attention recently (Levine et al., 2022; Chan et al., 2022; Shi et al., 2023b), but the problem is far from being solved. The central finding of this work is that structuring training data to increase semantic interdependence is an effective strategy towards better long context utilization. We achieve this by introducing and evaluating Structured Packing for Long Context (SPLiCe), a method for creating training examples by using retrieval (e.g., BM25, Contriver or repository structure) to collate mutually relevant documents into a single training context.
Figure 1: Training examples generated by the standard sampling procedure vs proposed by SPLiCe. Similar colors indicate related documents, which could be found using a retrieval method (e.g., BM25 or contriver) or metadata (e.g., repository structure for code).
@@
We empirically validate SPLiCe showing that fine-tuning of OpenLLaMA $3$Bv2, $7$Bv2 (Geng and Liu, 2023) and CodeLlama $13$B (Rozière et al., 2023) with mere $2$B–$6$B tokens already brings substantial improvements in handling long context information in downstream tasks that require retrieval and in-context learning. These tasks include Qasper (Dasigi et al., 2021) from SCROLLS (Shaham et al., 2022), lost-in-the-middle benchmark (Liu et al., 2023), HotPotQA (Yang et al., 2018), TREC (Li and Roth, 2002; Hovy et al., 2001), and DBpedia (Lehmann et al., 2015). We perform a comprehensive study of the design choices and properties of SPLiCe, showing, in particular, that the acquired long context capabilities transfer between modalities, such as code and text. SPLiCe also helps to retain and in some cases even improve performance on short context benchmarks like GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2021) and TREC (Li and Roth, 2002; Hovy et al., 2001).
@@
  • We comprehensively show that structuring training data is a viable way of improving the long context utilization of LLMs. To this end, we introduce SPLiCe, a method for creating training examples by using retrieval to collate mutually relevant documents into a single training context.
  • We fine-tune OpenLLaMA $3$Bv2, OpenLLaMA $7$Bv2 (Geng and Liu, 2023) and CodeLlama $13$B (Rozière et al., 2023) using SPLiCe, showing that it improves downstream performance on long context tasks.
  • We present a thorough analysis of the design choices behind SPLiCe, such as the number of retrieved documents and the order in which documents are merged into a training example.
@@ 2 Method
SPLiCe is a method for constructing training examples that enhance the LLMs’ long context utilization. This translates to improved performance on a range of tasks like in-context learning, question answering, in-context information retrieval, and long-context language modeling.
Algorithm 1 SPLiCe training example construction
Input:
   $D$: document corpus
   $k$: breadth hyper-parameter
   $L$: maximum length of returned training example
   RETRIEVE: retrieval method to use, e.g., BM25
   ORDER: ordering method, e.g., identity, or random shuffle
Output: training example consisting of concatenated documents
@@
return $\texttt{TRIM}(\texttt{ORDER}(C),L)$
Rationale and intuitions
Capturing long-range dependencies is believed to enhance language modeling and retrieval-augmentation (Borgeaud et al., 2022). However, it is an open question how to achieve such benefits in pre-training or fine-tuning. The primary difficulty comes from long-range dependencies being rare in training data (de Vries, 2023) and diminishing with distance. Thus, it is unlikely that a model will learn to utilize long context without more explicit guidance.
Recent studies indicate that structuring data, i.e., going beyond the i.i.d. paradigm, might be beneficial or even necessary to achieve good long-context performance. (Levine et al., 2022) develops a theory showing that the trained model establishes stronger dependencies between text segments in the same training example. Chan et al. (2022) indicates that specific distributional properties of the training data have a significant impact on the model’s in-context performance. Finally, concurrently to our work (Shi et al., 2023b) shows that pre-training on structured data improves the performance of LLMs.
SPLiCe follows these intuitions. Namely, it constructs training examples by collating mutually relevant documents to increase the dependency density, thus allowing the model to learn to utilize long context. We provide additional insights into the distributional properties of SPLiCe created data at the end of this section and in Appendix K.
Structured Packing for Long Context (SPLiCe)
SPLiCe starts by picking a random document from the dataset to create a root of a tree and continues in a breadth-first manner, each time appending top-$k$ similar documents from the corpus. The final sequence consists of the tree flattened according to a certain tree traversal strategy; see Algorithm 1.
The hyperparameter $k$ introduces flexibility, enabling interpolation between different modes of retrieval. Specifically, when $k=1$, SPLiCe simulates a long document by creating a path of related examples. For larger $k$ values, SPLiCe generates examples akin to those used in retrieval-augmented models, e.g., (Borgeaud et al., 2022). It is essential for the dataset to be well-curated to ensure the efficient functioning of SPLiCe. For instance, maintaining a low duplication rate encourages the model to leverage long context beyond merely copying. In this work, we employ the StarCoder (Li et al., 2023b) dataset, which has been deduplicated using the pipeline described in (Allal et al., 2023).
  • SPLiCe Repo : based on additional meta-information about the data, that is the repository structure of the code (Repo): we concatenate files using a depth-first search algorithm on the directory structure, that is files from the same directory are grouped together. A similar method has been pioneered by (Wu et al., 2022) and proposed in (Shi et al., 2023b) as an interesting future direction.
  • SPLiCe BM25 : based on BM25 (Robertson and Zaragoza, 2009; Bassani, 2023), a standard retrieval method that uses a bag-of-words approach to rank documents based on their similarity to a query.
  • SPLiCe Cont : based on Contriever-MSMARCO (Cont) (Izacard et al., 2022), a recently proposed retrieval method that uses a transformer to rank documents based on their similarity to a query.
Baselines One standard approach, commonly used in LLM training pipelines, is to randomly sample documents from the dataset and concatenate them to make training examples of a desired context. This is known as example packing (Brown et al., 2020). Another popular approach restricts sampled documents to the same meta-class (for example, Wikipedia article, source code) (Groeneveld et al., 2024; Zhao et al., 2024). We compare SPLiCe against both.
SPLiCe computational efficiency Given the dataset sizes used in training LLMs, computational efficiency plays a non-trivial role. SPLiCe Repo is the fastest and easiest to implement but requires additional directory structure, i.e., it does not apply to general web data. SPLiCe BM25 uses a bag of words BM25 method that lacks deeper semantic encoding. SPLiCe Cont requires calculating embeddings for each document and retrieval based on the vector inner-product. The latter step can be done efficiently using fast approximate inner-product search, e.g., Faiss (Johnson et al., 2017).
Burstiness SPLiCe conjecturally falls into the framework presented in Chan et al. (2022), which shows that the distributional properties of the training data affect the in-context capabilities of transformer models. In particular, it indicates the importance of "burstiness", i.e., a flatter frequency distribution with a relatively higher mass on the rare, long-tail tokens appearing in a sequence. In Appendix K, we show that SPLiCe increases the burstiness of the training data (measured in terms of Zipf’s coefficient of token frequency) compared to the example packing baseline.
@@ 3 Experiments
In this section, we show that SPLiCe improves the long context performance of large-scale language models. To this end, we use $3$B, $7$B, and $13$B parameter models and evaluate on tasks that test in-context learning (Section 3.2.1), question answering (Section 3.2.2), and in-context information retrieval capabilities (Section 3.2.3). Additionally, we provide perplexity results in Appendix E. Next, we show that SPLiCe can improve the core model capabilities by testing the short context performance, see Section 3.2.5. Finally, in Section 3.3, we tune over 30 medium ($270$M parameters) models using different data mixtures and SPLiCe parameters, to analyze its design choices.
An important finding of our work is that presented improvements occur during a relatively short, compared to pre-training, fine-tuning. To be more precise, $3$B models were tuned on $5.4$B tokens, whereas $7$B and $13$B models were tuned on $2$B tokens. The $3$B and $7$B start from OpenLLaMAv2 (Geng and Liu, 2023) that was pre-trained on $1$T tokens, the $13$B parameter models start from CodeLlama (Rozière et al., 2023) that was pre-trained on $2.5$T tokens.
@@ 3.1 Experimental setup
For $3$B experiments, we continue training on a 50/50 mixture of RedPajama, prepared in the standard way, and C prepared using SPLiCe BM25 . For $7$B and $13$B ones, we continue the training on a 50/25/25 mixture of RedPajama (50) prepared in the standard way, StackExchange (25) and C (25) prepared using SPLiCe BM25 . StackExchange is part of RedPajama TogetherComputer (2023), and C data come from StarCoder (Li et al., 2023b). Including the standard RedPajama aims to prevent the model from overfitting to artificially created long documents. We refer to Appendix A for a detailed description of the model architectures.
We train with $32$K context length. We use a batch size of $256$K tokens per step and the learning rate of ${1.5}\mathrm{e}{-5}$ with linear warmup and cosine decay, following (Geng and Liu, 2023). For $13$B model we use batch of $512$K tokens per step.
Regarding the context extensions method, we use FoT (Tworkowski et al., 2023), and the CodeLlama (CL) (Rozière et al., 2023) approach. More explicitly, we test eight models:
@@
where Baseline denotes the standard (example packing) data preparation method where context is created by sampling random documents from the corpus and separating them with BOS/EOS tokens. We provide results regarding the within-domain sampling baseline in Appendix C.2. If not stated otherwise, in SPLiCe we use $k=1$ and the identity permutation as Order in the Algorithm 1. Hyperparameter details can be found in Appendix A.
@@ 3.2 Experimental results
3.2.1 In-Context learning
To evaluate the improvements of in-context learning abilities, we use TREC (Li and Roth, 2002; Hovy et al., 2001) and DBpedia (Lehmann et al., 2015), which are text classification tasks. For TREC, to test the long context utilization, we vary the number of in-context examples in $\{780,1560\}$ (which correspond roughly to $\{16,32\}$K context length). We sample these examples multiple times and measure classification accuracy on the test set. In Table 1, we observe that SPLiCe improves the average performance across different context lengths and model sizes. We follow a very similar protocol for DBpedia ($\{16,32\}$K context length corresponding to $\{190,380\}$ in-context examples), confirming that SPLiCe improves the context utilization. In Appendix I, we study the distribution of improvements with respect to the choice of the in-context examples, showing the stochastic domination of SPLiCe over the baseline.
Table 1: We test the average classification performance on TREC (Li and Roth, 2002; Hovy et al., 2001). We average across $50$ sets of in-context examples for $3$B models, $10$ sets for $7$B models, and $5$ sets for $13$B models. We use $\Delta$ [confidence interval] to denote the mean improvement and its $95\%$ bootstrap confidence intervals (see Appendix I). For DBpedia, we average results across $40$ sets of in-context examples for $3$B and $7$B models and $5$ for $13$B ones. Due to the size of the DBpedia evaluation dataset, for each set of in-context examples, we subsample a subset of $500$ elements of the evaluation set.
Task TREC DBpedia
Model Context Baseline SPLiCe $\Delta$ [conf interv] Baseline SPLiCe $\Delta$ [conf interv]
@@ 3.2.1 In-Context learning
$16$K 88.2 91.2 3.0 [1.9,3.5]{1.9,3.5]}1.9 , 3.5 ] 95.8 96.8 1.0 [0.6,1.5]{0.6,1.5]}0.6 , 1.5 ]
Table 2: We measure question answering over long input using Qasper (Dasigi et al., 2021) (2-shot setting) and HotPotQA (Yang et al., 2018) (10-shot setting). For Qasper, we use the implementation from Language Model Evaluation Harness (Gao et al., 2021). For HotPotQA, we average results across $7$ sets of in-context examples for $3$B and $7$B models and $2$ for $13$B ones, with $\Delta$[confidence interval] denoting mean improvement and its $95\%$ bootstrap confidence intervals. Note that in the $3$B model case, despite using SPLiCe for code data only, we still have improvements in non-code tasks.
Task QASPER HotPotQA
Model Context Baseline SPLiCe Context Baseline SPLiCe $\Delta$ [conf interv]
@@
3.2.2 Question answering
We also show that SPLiCe models have improved question answering capabilities. To this end, we use Qasper (Dasigi et al., 2021) from SCROLLS (Shaham et al., 2022) and HotPotQA (Yang et al., 2018). Qasper contains questions regarding research papers provided as a part of the input context. In Table 2, we observe that SPLiCe improves the two-shot performance. Similarly, we observe improvements on HotPotQA, which tests the ability to aggregate knowledge across multiple documents. We stress that those results measure few-shot performance, as neither of the datasets was used during training.
3.2.3 Lost in the middle
The key retrieval task introduced in (Liu et al., 2023) has become a routine diagnostic tool to study context capabilities. Specifically, the model is presented with a list of key-value pairs, which are $128$-bit UUIDs, and is tasked to retrieve the value for a given key. Even though very simple, the task proved to be challenging for many open-source language models. Specifically, LMs tend to achieve better performance when the relevant information is either in the beginning or the end of the input context, while struggling in the middle. See Appendix D for the task details.
In Figure 2, we present the key retrieval performance of $7$B models trained with CodeLlama (Rozière et al., 2023) context extension method. We note that despite relatively short tuning, SPLiCe significantly helps on hard-to-retrieve positions (the position of the key in the range $[25,100]$). We present more details in Appendix D.
Figure 2: Key-value retrieval performance on a dictionary of $300$ key-value pairs ($\approx$ $24$K tokens). The $7$B CL model trained with SPLiCe achieves much higher accuracy on hard-to-retrieve positions in the middle. The details about this task can be found in Appendix D. Each position was evaluated using $500$ examples.
3.2.4 Ordering of examples
We typically use the identity ordering as Order in Algorithm 1 to merge documents into a single context, as we found that in most cases, it performs best. In Appendix F, we found that random shuffling is slightly better in some cases. Specifically, this is the case of large $7$B CodeLlama models, see Table Table 18. We hypothesize that random ordering forces the model to make use of the full space of the RoPe positional encoding. Whereas the identity ordering, also used in (Shi et al., 2023b), skews the model toward paying more attention to fragments of text that are not too far away. This suggests an interesting research direction on the intersection of data preparation and positional embeddings.
3.2.5 Short context evaluation
An undesirable side effect of fine-tuning for extended context (e.g., with SPLiCe) could be performance degradation on shorter contexts. To test this, we evaluate the short context performance of our models using GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2021) and TREC (Li and Roth, 2002; Hovy et al., 2001). We present the results in Table 3. We note that for smaller models, both SPLiCe and Baseline show some short context degradation, whereas the larger $13$B CodeLlama Model SPLiCe brings significant improvements in short context performance on GSM8K ($+1.7$), TREC ($+3.3$) and maintains the MMLU performance. We provide an additional evaluation on code in Appendix L.
Table 3: We evaluate our models on MMLU ($5$-shot), GSM8K ($8$-shot) and TREC ($90$-shot). We provide an additional comparison with their starting checkpoint. For $7$B case, we additionally compare with a model tuned with $2$k context length on the same data. For each task, we highlight best results up to $1$ point.
Model MMLU GSM8K TREC
STEM Humanities Social Other Mean
$3$B Starting Chkpt 24.4 25.6 26.3 27.8 25.9 4.3 58.4
@@ 3.2.5 Short context evaluation
$7$B Starting Chkpt 33.6 42.1 46.0 44.5 40.8 8.0 59.6
$7$B 2K Tuned 33.7 40.8 44.6 41.1 39.4 8.4 62.3
[…]
$13$B Starting Chkpt 36.6 48.2 50.6 45.4 44.3 21.4 78.7
[…]

### v9

@@ Abstract
Recent advancements in long-context language modeling have attracted significant attention, yet their practical applications often suffer from suboptimal context utilization. To efficiently address this issue, we introduce the Structured Packing for Long Context, SPLiCe, a method that uses retrieval to collate mutually relevant documents into long training samples. We demonstrate that SPLiCe improves performance on long-context tasks, particularly by achieving perfect accuracy on the synthetic Needle in the Haystack benchmark, and effectively mitigating the ‘lost-in-the-middle’ phenomenon often observed in large language models. Notably, these long-context capabilities also extend to realistic downstream tasks, such as Qasper, across multiple model sizes—3B, 7B, and 13B—and are achieved with only brief fine-tuning on 2-6 billion tokens. We supplement these results with a detailed analysis of SPLiCe, examining the impact of hyperparameter choices, the different mixtures and proportions of SPLiCe-generated training data, and the choice of the retriever. We also study the transfer of long-context utilization skills between the modalities. An intriguing finding from our analysis is that training on a corpus of code can enhance performance on natural language tasks.
Codehttps://github.com/ideas-ncbr/publications˙2024
Extended versionhttps://arxiv.org/abs/2312.17296
@@ 1 Introduction
Large language models (LLMs) (Brown et al. 2020; Chowdhery et al. 2022; Lewkowycz et al. 2022; OpenAI 2023a; Bai et al. 2023) have transformed the field of AI. Recently, the field has observed the rise of Long Context Language Models (LCLMs) that promise to unveil novel and powerful capabilities (Anthropic 2023; OpenAI 2023b; Gemini Team 2024). However, their ability to process long contexts is not always as effective as one hopes. Indeed, several studies have highlighted an important limitation: when processing prompts composed of multiple documents, LCLMs frequently encounter difficulties in accurately extracting relevant information (Tworkowski et al. 2023; Liu et al. 2023; Shi et al. 2023; Kamradt 2023). Additionally, they typically find it challenging to utilize information from the middle of their inputs (Liu et al. 2023), even on simple synthetic retrieval tasks (Li et al. 2023a). Understanding these issues is vital for advancements in LCLM technologies and calls for systematic research.
@@
In this work, we take a step towards better context utilization in LCLMs. We focus on training data, keeping other components, such as the architecture and training objectives, unchanged. The broad question is: Given training data consisting of documents, how should these documents be organized into training samples to enhance long-context capabilities? While this perspective has received some attention recently (Levine et al. 2022; Chan et al. 2022; Shi et al. 2024), the problem remains unsolved. The central finding of this work is that structuring training data to increase semantic interdependence is an effective strategy towards better long context utilization. We achieve this by introducing and evaluating Structured Packing for Long Context (SPLiCe), a method for creating training samples by using retrieval (e.g., BM25, Contriever) to collate mutually relevant documents into a single training context.
@@
We empirically validate SPLiCe showing that fine-tuning of OpenLLaMA $3$Bv2, $7$Bv2 (Geng and Liu 2023) and CodeLlama $13$B (Rozière et al. 2023) with mere $2$B–$6$B tokens already brings improvements in handling long context information in downstream tasks that require retrieval and in-context learning. These tasks include Qasper (Dasigi et al. 2021) from SCROLLS (Shaham et al. 2022), HotPotQA (Yang et al. 2018), Needle In A Haystack (Kamradt 2023), TREC (Li and Roth 2002; Hovy et al. 2001), and DBpedia (Lehmann et al. 2015). We also show that SPLiCe significantly alleviates the ’lost-in-the-middle’ phenomenon (Liu et al. 2023) and outperforms standard example packing on the Needle In A Haystack task (Kamradt 2023) (see Figure 1). We perform a comprehensive study of the design choices and properties of SPLiCe, showing, in particular, that the acquired long context capabilities transfer between modalities, such as code and text. SPLiCe also helps to retain and in some cases even improve performance on short context benchmarks like GSM8K (Cobbe et al. 2021), MMLU (Hendrycks et al. 2021) and TREC (Li and Roth 2002; Hovy et al. 2001).
Figure 1: SPLiCe vs Example Packing (EP) (baseline) on Needle in a Haystack. A model fine-tuned with SPLiCe achieves perfect accuracy in retrieving fine-grained information over the whole context, while the baseline can only handle a small final segment (details in Appendix N).
@@
  • We comprehensively show that structuring training data is a viable way of improving the long context utilization. To this end, we introduce SPLiCe, a method for creating training samples by using retrieval to collate mutually relevant documents into a single sample.
  • We fine-tune OpenLLaMA $3$Bv2, OpenLLaMA $7$Bv2 (Geng and Liu 2023) and CodeLlama $13$B (Rozière et al. 2023) using SPLiCe, showing that it improves long-contex downstream performance.
  • We provide a comprehensive analysis of SPLiCe’s design choices, including retrieval parameters and document concatenation order, and evaluate its robustness and scalability with varying data sources and a parametrizable noisy retriever.
@@ 2 Method
SPLiCe is a method for constructing training samples that improve the effectiveness of long-context fine-tuning. This leads to improved performance in tasks such as in-context learning, question answering, information retrieval, and long-context language modeling (see Section 3).
Rationale and Intuitions
Capturing long-range dependencies is believed to enhance language modeling and retrieval-augmentation (Borgeaud et al. 2022). It is an open question how to achieve such benefits in pre-training or fine-tuning. The primary difficulty comes from long-range dependencies being rare in training data (de Vries 2023) and diminishing with distance. Thus, it is unlikely that a model will learn to utilize long context without more guidance.
Recent studies indicate that structuring data, i.e., going beyond the i.i.d. paradigm, might be beneficial or even necessary to achieve good long-context performance. (Levine et al. 2022) develops a theory showing that the trained model establishes stronger dependencies between text segments in the same training sample. Whereas concurrently to our work (Shi et al. 2024) shows that pre-training on structured data improves performance (see also Section 4 for a more detailed comparison). SPLiCe follows these intuitions, and constructs training samples by concatenating mutually relevant documents to increase the dependency density, thus allowing the model to learn to utilize long context.
Structured Packing for Long Context (SPLiCe)
SPLiCe starts by picking a random document from the dataset to create a root of a tree and continues in a breadth-first manner, each time appending top-$k$ similar documents from the corpus. The final sequence is generated by flattening the tree according to a specific traversal strategy; see Algorithm 1. The hyperparameter $k$ introduces flexibility, enabling interpolation between different retrieval modes. Specifically, when $k=1$, SPLiCe simulates a long document by creating a path of related examples. For larger $k$ values, SPLiCe generates examples akin to those used in retrieval-augmented models, e.g. (Borgeaud et al. 2022).
SPLiCe Retrieval
  • SPLiCeRepo: based on additional meta-information about the data, that is the repository structure of the code (Repo): we concatenate files using a depth-first search algorithm on the directory structure, that is files from the same directory are grouped together. A similar method has been pioneered by (Wu et al. 2022) and proposed in (Shi et al. 2024) as an interesting future direction.
  • SPLiCeBM25: based on BM25 (Robertson and Zaragoza 2009; Bassani 2023), a standard retrieval method that uses a bag-of-words approach to rank documents based on their similarity to a query.
  • SPLiCeCont: based on Contriever-MSMARCO (Cont) (Izacard et al. 2022), a retrieval method that uses a transformer to rank documents based on their similarity.
Figure 2: Training samples generated by Example Packing, Within-Domain Example packing, and SPLiCe. Similar colors and shapes indicate related documents, which could be found using a retrieval method (e.g., BM25 or Contriever) or metadata (e.g., git repository structure).
Figure 3: Key-value retrieval performance on a dictionary of $300$ key-value pairs ($\approx$ $24$K tokens). The $7$B CL model trained with SPLiCe achieves much higher accuracy on hard-to-retrieve positions in the middle than the Example Packing Baseline. The details about this task can be found in Appendix D. Each position averaged over $500$ examples.
Algorithm 1 SPLiCe training sample construction
  Input:    $D$: document corpus   $k$: breadth hyper-parameter   $L$: maximum length of returned training sample   RETRIEVE: retrieval method to use, e.g., BM25   ORDER: ordering method, e.g., identity, or shuffle
  Output: training sample
@@
  return $\texttt{CONCAT}(\texttt{TRIM}(\texttt{ORDER}(C),L))$
SPLiCe Computational Efficiency
Given the dataset sizes used in training LLMs, computational efficiency plays a crucial role. SPLiCeRepo is the fastest and easiest to implement but requires additional directory structure, i.e., it does not apply to general web data. SPLiCeBM25 uses a bag of words BM25 method that lacks deeper semantic encoding. However, it was observed to have strong generalization properties (Thakur et al. 2021). SPLiCeCont requires calculating embeddings for each document and retrieval based on the vector inner-product, but can have poorer generalization properties than BM25 (Thakur et al. 2021). The retrieval step can be done efficiently using a fast approximate max IP search, e.g., Faiss (Johnson, Douze, and Jégou 2017). To reduce the number of times the training sample requires just copy-paste abilities and improve training step efficiency, we employ the StarCoder (Li et al. 2023b) dataset, which was deduplicated using the pipeline from (Allal et al. 2023).
@@ 3 Experiments
In this section, we show that SPLiCe improves the long context performance of large-scale language models. To this end, we use $3$B, $7$B, and $13$B parameter models. First, in Section 3.3, we focus on tasks that test in-context learning, question answering, and in-context information retrieval. Next, we show that SPLiCe can improve the core model capabilities by testing its short context performance. Finally, in Section 3.4, we train over 40 medium-size models ($270$M parameters) using different data mixtures and SPLiCe parameters to analyze various design choices, robustness to noise, and scaling properties.
An important finding of our work is that presented improvements occur during a relatively short fine-tuning. To be more precise, $3$B models were tuned on $5.4$B tokens, whereas $7$B and $13$B models were tuned on $2$B tokens.
3.1 Baselines
We consider two popular baselines used in LLM training pipelines. The first one is Example Packing (Brown et al. 2020), used in the training of GPT-3 models. It constructs training samples by randomly sampling documents from the dataset and separating them with BOS/EOS tokens. The second one, which we call Within-Domain Example Packing takes random documents from the same meta-class (for example, Wikipedia, C source code) and concatenates them to create a training sample (Groeneveld et al. 2024; Zhao et al. 2024). We compare SPLiCe against both baselines. As we note no clear benefit of Within-Domain Example Packing over Example Packing in fine-tuning case (see Table 21 in Appendix B.3 ) in the main body of the paper we compare only against a more established Example Packing. We visualize the differences between baselines in Figure 2.
@@ 3.2 Experimental Setup
For $3$B model experiments, we fine-tune on a 50/50 mixture of RedPajama, prepared in the standard way, and C prepared using SPLiCeBM25. For $7$B and $13$B ones, we fine-tune on a 50/25/25 mixture of RedPajama (50) prepared in the standard way, StackExchange (25) and C (25) prepared using SPLiCeBM25. StackExchange is part of RedPajama (TogetherComputer 2023), and C data come from StarCoder (Li et al. 2023b). Including the standard RedPajama aims to prevent the model from overfitting to artificially created documents and is inspired by (Ouyang et al. 2022; Rozière et al. 2023). We analyze the impact of data mixture in Section 3.4.
We fine-tune with $32$K context length. We employ the Focused Transformer (FoT) (Tworkowski et al. 2023) and CodeLlama (CL) context extension methods (Rozière et al. 2023). We use a batch size of $256$K ($512$K, resp.) tokens per step for $3$B and $7$B ($13$B, resp.) models. We set the learning rate of ${1.5}\mathrm{e}{-5}$ with linear warmup and cosine decay, following (Geng and Liu 2023). In the next section, we test eight models:
@@
where EP denotes the standard Example Packing (Brown et al. 2020) method (serving as baseline) where context is created by sampling random documents from the corpus and separating them with BOS/EOS tokens (see Figure 2). We provide results regarding the Within-Domain Example Packing baseline in Appendix C.2. If not stated otherwise, in SPLiCe we use $k=1$ and the identity permutation as Order in the Algorithm 1. Hyperparameter details can be found in Appendix A.
@@ 3.3 Experimental Results
In-Context Learning
In this section, we ask the following research questions: Does SPLiCe improve in-context learning abilities? If so, with what context length is it the case? To answer those questions, we evaluate the accuracy of our models on TREC (Li and Roth 2002; Hovy et al. 2001) and DBpedia (Lehmann et al. 2015), which are text classification tasks. For TREC we test $\{2,16,32\}$K context lengths, which correspond to $\{90,780,1560\}$ in context examples, respectively. For DBpedia, we test $\{16,32\}$K context lengths, which correspond to $\{190,380\}$ in-context examples, respectively, and omit the 2K length due to its limited capacity of 20 in-context examples. For each context length, we average the results across several sets of in-context examples and provide average improvement of SPLiCe and its $95$% bootstrap confidence interval (improvements are calculated per set of in-context examples, see Appendix H). In both tasks and all considered context lengths, we note that SPLiCe significantly improves in-context learning abilities in comparison to both Example Packing and the starting checkpoint. We hypothesize that by increasing the amount of potentially relevant information in context, SPLiCe allows the model to learn longer and better context lookups. We further study this in Section 3.5, where we analyze SPLiCe using the framework from (Chan et al. 2022). We present the main results in Tables 1, 2 and additional in Table 35 in Appendix O. Additionally in Appendix H we analyze results per-set of in-context examples and show that SPLiCe achieves stochastic domination over Example Packing.
TREC
Model Context EP SPLiCe $\Delta$ [conf interv]
@@
$16$K 88.2 91.2 3.0 [1.9,3.5]{1.9,3.5}]1.9 , 3.5 ]
Table 1: We test the classification accuracy on TREC (Li and Roth 2002; Hovy et al. 2001). We average across $50$ sets of in-context examples for $3$B models, $10$ for $7$B models, and $5$ for $13$B models. $\Delta$ [ci] denotes the mean improvement and its $95\%$ bootstrap confidence intervals (see Appendix H).
DBpedia
Model Context EP SPLiCe $\Delta$ [conf interv]
[…]
$16$K 79.1 82.0 2.9 [2.5,3.4]{2.5,3.4}]2.5 , 3.4 ]
[…]
$16$K 83.6 85.6 2.0 [1.5,2.5]{1.5,2.5}]1.5 , 2.5 ]
[…]
$16$K 96.2 96.4 0.2 [0.0,0.3]{0.0,0.3}]0.0 , 0.3 ]
[…]
$16$K 95.8 96.8 1.0 [0.6,1.5]{0.6,1.5}]0.6 , 1.5 ]
Table 2: We average results across $40$ sets of in-context examples for $3$B and $7$B models and $5$ for $13$B. Due to the size of the DBpedia dataset, for each set of in-context examples, we sample a subset of $500$ elements of the evaluation set.
Question Answering and In-Context Retrieval
In this section, we ask the following research question: Does fine-tuning on SPLiCe prepared data result in improved question-answering abilities? To answer the question, we utilize popular long context benchmarks such as Needle In A Haystack (Kamradt 2023) and lost-in-the-middle key-value retrieval (Liu et al. 2023), along with Qasper (Dasigi et al. 2021) from SCROLLS (Shaham et al. 2022), HotPotQA (Yang et al. 2018) passkey (Mohtashami and Jaggi 2023) and parts of RULER (Hsieh et al. 2024) tasks.
On Needle In A Haystack, we observe that the model fine-tuned on data prepared by SPLiCe strongly outperforms the model fine-tuned on data prepared by Example Packing. To be more precise model fine-tuned with SPLiCe can answer the question no matter the location of the relevant piece of information. Whereas the model trained with Example Packing only manages to answer correctly when the information is close to the question (see Figure 1). We also test our models on the lost-in-the-middle key-value retrieval task (Liu et al. 2023), and observe that SPLiCe helps on hard-to-retrieve positions (see Figure 3). The main difference between those two tasks is that in the lost-in-the-middle key-value retrieval task, the input is highly structured (dictionary of random $128$ bit UUIDs, see Appendix D for details) and the objective of the model is to retrieve the value assigned to a given key. On the other hand, in the Needle In A Haystack, a piece of information is placed inside a large coherent text, and the model is asked a question related to this information (see Appendix N for details).
We additionally evaluate our models on Qasper (Dasigi et al. 2021), HotPotQA (Yang et al. 2018) passkey retrieval (Mohtashami and Jaggi 2023) and RULER (Hsieh et al. 2024) and observe that SPLiCe results in improvements over both the Example Packing and starting checkpoint. We present the results in Appendix K.
Short Context Evaluation
One challenge in long-context fine-tuning is the degradation of short-context performance (Dubey, Jauhri, and et al. 2024). This can be overcomed by upsampling the short-context data and more gradual context extension (Dubey, Jauhri, and et al. 2024). We note that those approaches are compatible with SPLiCe and instead focus on comparing SPLiCe with Example Packing in a single-step context extension setup. We observe that SPLiCe seems to be either better or on par with Example Packing (see Table 3). What is intriguing is that for the $13$B parameter model, SPLiCe even improves the short context performance on GSM8K (Cobbe et al. 2021) by ($+1.7$) over the starting checkpoint. We hypothesize that GSM8K is a much more attention-demanding task than MMLU (Hendrycks et al. 2021), as it requires extracting relevant pieces of information, composing a chain of thought, and writing the final answer. Whereas MMLU is a well-established collection of tests spanning multiple domains. We hypothesize that the improvement does not occur in smaller models due to their low scores on GSM8K, as we get similar results when evaluating on code in Appendix L.
Model MMLU GSM8K
STEM HUM All
$7$B ST CHKP 33.6 42.1 40.8 8.0
$7$B 2K Tuned 33.7 40.8 39.4 8.4
@@
$13$B ST CHKP 36.6 48.2 44.3 21.4
@@
Table 3: We evaluate our models on MMLU ($5$-shot), GSM8K ($8$-shot CoT). We provide an additional comparison with their starting checkpoint. For the $7$B case, we additionally compare with a model tuned with $2$k context length on the same data. For each task, we highlight the best results up to $1$ point. For $3$B model results see Appendix I.